he peak inspiratory pressue is 38 cmlH20 and the PEEP is 2 cmH20 with a set Vi of 900 mL. The tubing compliance factor is 3.0 mL/cmH20 a. what is the lost volume: b. delivered volume: e peak inspiratory pressue is 50 cmH20 and the PEEP is O cmI120 with a set Vt of 500 mL. The tubing compliance factor is 3.6 mL/cmH20 a. what is the lost volume: b. delivered volume;
Answer #1

Volume lost= compliance factor*peak pressure that is 3ml/ cmH2o *38cm H2o= 114. And what is delivered to patient is Vt - volume lost=900-114=786ml, 2) volume lost=180 cm H2O , what is delivered to patient is 320 ml